OpenStudy (anonymous):

You want to double your savings of $29,952 in an account with an interest rate of 2.76% that compounds daily. How long will it take you to double your money? Round the solution to the nearest hundredth. Well I'm stuck in this part: 2 = 1.00007561643836^365t *I used the formula: A=P(1+r/n)^nt

OpenStudy (anonymous):

I don't really want to explain it because it would take like 30 minutes but I can link you to a yahoo answers if you want?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

I sent it to you in a pm

OpenStudy (anonymous):

like i jsut dont know what "log'' means... (i think that's what u gotta use right?

OpenStudy (anonymous):

the yahoo link... doesnt explain it... or at least not how i did it... like, i dont get it...

OpenStudy (anonymous):

Do you have your text book with you? If so it should show how to do it in the front of the lessons

OpenStudy (anonymous):

online class... it doesnt... (Im suppose to know it... ) but last year i was a mess, and now i wanna change...

OpenStudy (anonymous):

how to get the "log" of that... i jsut dont get what a "log" is

OpenStudy (anonymous):

I go to Connections Academy and I have text books, the only log I know is a logarithm which I don't think has much to do with this lol

OpenStudy (anonymous):

But anyways I've got to go, my mom needs me to put her table together... goodluck

OpenStudy (ryan123345):

look up logarithium

OpenStudy (anonymous):

A = P (1+r/n)^nt 59905 = 29952(1+.0276/365)^365t 2 = 1.00007561643836^365t stuck here! :/

OpenStudy (anonymous):

nobody??????